OpenStudy (anonymous):

HELP!!!! Evaluate the integral in Polar Coordinates. R Is bounded by y=sqrt(4x-x^2), y=x. The double integral is...

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\int\limits_{R}^{?}\sqrt{x ^{2}+y ^{2}}dA\]

OpenStudy (anonymous):



OpenStudy (anonymous):

How do I find the limits for dr and dtheta?

OpenStudy (anonymous):

Some preliminary work on my part: \[\sqrt{4x-x^2}=\sqrt{-\left(x^2-4x+4-4\right)}=\sqrt{4-(x-2)^2}\] Region: \[\displaystyle\int\int_R\sqrt{x^2+y^2}~dA\]